OpenStudy (anonymous):

9a^2+24ab+16b^2 Factoring

mathslover (mathslover):

Try to write the given expression in the form of \(a^2 - 2ab + b^2\)

OpenStudy (anonymous):

It is

mathslover (mathslover):

Not perfectly. We need to simplify it more. Like, writing 9a^2 as (3a)^2 ...

OpenStudy (anonymous):

So 3a^2+8ab+4a^2?

mathslover (mathslover):

No, first of all, don't forget to put brackets : \((3a^2) + 8ab + (4b)^2 \) Now, the second term is wrong, we had 24ab and you have written 8ab, we can write it as : \(24ab = 2(3a)(4b)\) \(\implies (3a)^2 + 2(3a)(4b) + (4b)^2\)

OpenStudy (anonymous):

Wait so the bottom equation is the answer?

mathslover (mathslover):

Now, compare it with : \(a^2 + 2ab + b^2\) You get , a = 3a, b = 4b... Since, \(a^2+ 2ab + b^2 = (a+b)^2\) Therefore, \((3a)^2 + 2(3a)(4b) + ( 4b)^2 = (3a+4b)^2 \textbf{or} ~ (3a+4b)(3a+4b)\)

mathslover (mathslover):

That is not the answer, it was yet to be factored ...

OpenStudy (anonymous):

Ah this why i failed the class the first time. lol. So from (3a+4b)(3a+4b)

mathslover (mathslover):

Yes that is the answer now.